The Cycle More Company manufactures high end bicycles. The controller for the company Ima Sewsique, spent much of the past qarter at home due to a lingering illness. As a result the company records are incomplete. The existing records are presented below Beginning Balances 1/1/2018 Raw Materials Inventory $7,500 $7,500 Finished Goods Inventory Ending Balances 12/31/2018 Raw Materials Inventory Work in Process Finished Goods Inventory $1,525 $1,025 Additional Information: $98,000 $305,000 $103,000 $143,000 400% Raw Materials Purchased Raw Materials Used in Production Prime Costs Overhead Rate based on Direct Labor Dollars Required: 1. Prepare T-accounts for all of the Manufacturing Inventory Accounts and Cost of Goods Sold. Determine Conversion costs for the period and the total manufacturing costs added to the Work In Process. List any direct costs and explain how your determination was made. 2. 3. Consider the composition of manufacturing overhead. Label the items listed as either a variable cost, a fixed cost, or a mixed cost. 4. The owners of Cycle More have chosed to apply overhead based on direct labor dollars. Given the composition of your list of overhead items in your response to question 3 (above), when would direct labor be an accurate predictor of the cost of overhead? 5.The Cycle More Company manufactures high end bicycles. The controller for the company Ima Sewsique, spent much of the past quarter at home due to a lingering illness. As a result the company records are incomplete.
Answer #1

Working Note:

COST OF GOODS MANU. & COGS: Amount $
Ope WIP 7500
RM:
OB 6525
PUR 98000
LESS CB -1525
RM USED 103000
DL 40000
PRIME COST 143000
OVERHEAD 160000
MANUF.COST INCURRED 303000
TOTAL MANUFACTURING 310500
LESS WIP CB -1025  
COST OF GOODS MANUFACTURED 309475
FG OB 7500
FG AVAILABLE 316975
LESS: FG CB 11975
COGS 305000
1)
WIP INVENTORY A/c
Debit amount $ Credit amount $
OB 7500 FG 309475
DM 103000 CB 1025
DL 40000
OVERHEAD 160000
TOTAL 310500 TOTAL 310500
RAW MATERIAL A/c
OB 6525 WIP 103000
PURCH. 98000 CB 1525
TOTAL 104525 TOTAL 104525
FG INVENTORY A/c
OB 7500 COGS 305000
WIP 309475 CB 11975
TOTAL 316975 TOTAL 316975
COGS A/c:
FG INVEN. 305000
2)
Conversion costs = DL + overhead = 40000 + (40000*400%) = $200000
Total Manufacturing costs added to WIP = DM + CONVER. COSTS = 103000+200000 = $303000
3)
Direct Costs = DM + DL + OVERHEAD = $303000
Direct costs are production costs recoginsed at the factory level to Finish Goods.
4) Composition of manufacturing overhead:
indirect material variable
indirect labor variable
depreciation fixed
utilities semi-variable or mixed costs
5) The direct labor would be a accurate predictor of cost of overhead when
the production is a labor intensive manufacturing where labor constitutes
major portion of Production costs.